OpenStudy (anonymous):

Solve the following system using the substitution method. 3x – y = –2 6x + 5y = 24

OpenStudy (anonymous):

3x – y = –2 y=3x+2 substitute for y in , 6x + 5y = 24 6x + 5(3x+2) = 24 6x+15x+10=24 21x=14 x=14/21=2/3 put x back to get y y=3x+2=(3*2/3)+2=2+2=4 x=2/3, y=4